### v1

@@ 2.7 Comparing model-based and data-based SE(3) equivariance strategies
Model Temperature ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra $\tau$=1.2 15.45% 1.4224 0.7739 0.8348 0.9197 0.9468 0.9604 $\tau$=1.0 16.01% 1.3552 0.7837 0.8421 0.9227 0.9481 0.9618 $\tau$=0.8 16.30% 1.3643 0.7848 0.8429 0.9235 0.9491 0.9621 $\tau$=0.6 16.05% 1.4009 0.7794 0.8387 0.9220 0.9480 0.9627 $\tau$=0.4 15.86% 1.5081 0.7629 0.8261 0.9120 0.9442 0.9577 $\tau$=0.2 14.56% 1.9842 0.7005 0.7774 0.8743 0.9297 0.9323
@@ 2.8 Sampling temperature balances diversity and accuracy
Since a given molecular spectrum corresponds to a unique molecular structure, the task of molecular structure elucidation from spectra requires high accuracy in the generated results. However, diffusion models inherently introduce stochasticity during sampling due to the injection of random noise, which is essential for promoting diversity in the generated results. To better understand the trade-off between stochasticity and accuracy, we investigate how controlling the level of stochasticity affects performance. Specifically, we introduce a sampling temperature parameter $\tau$ in Eq. 6, which modulates the scale of the injected noise during the sampling process. Lower $\tau$ values reduce the influence of stochasticity, resulting in more deterministic outputs, while higher $\tau$ allows for greater exploration of the molecular space.
@@
As shown in Table 5, we evaluate DiffSpectra under varying $\tau$ values ranging from 0.2 to 1.2. We observe that moderate temperature values ($\tau=0.8$ and $\tau=1.0$) achieve the best balance between diversity and correctness. In particular, $\tau=0.8$ obtains the highest top-1 accuracy of 16.30%, while maintaining competitive similarity metrics, including a TanimotoMG of 0.7848, CosineMG of 0.8429, and TanimotoMA of 0.9235. Extremely low temperatures (e.g., $\tau=0.2$) result in overly deterministic generations with reduced diversity and higher MCES (1.9842), while extremely high temperatures (e.g., $\tau=1.2$) also degrade performance due to excessive stochasticity.
@@
Overall, these results suggest that selecting a moderate $\tau$ helps DiffSpectra strike an effective trade-off between generating structurally diverse candidates and maintaining high similarity to the ground-truth molecules.
3 Conclusion
In this work, we present DiffSpectra, a novel spectrum-conditioned diffusion framework for molecular structure elucidation. By integrating a continuous-time variational diffusion model with a Transformer-based molecular graph backbone, DiffSpectra is capable of generating molecular graphs with consistent 2D and 3D structures under spectral conditions. To encode spectral conditions, we introduce SpecFormer, a multi-spectrum Transformer encoder, which is further pre-trained via masked patch reconstruction and contrastive learning to align spectra with molecular structures. Through extensive experiments, we demonstrate that DiffSpectra achieves superior performance in recovering molecular structures from spectra, benefiting from the incorporation of pre-trained SpecFormer and multi-modal spectral inputs. In addition, our analysis shows that a model-based SE(3) equivariant architecture provides robust geometric consistency, while proper sampling temperature helps balance stochasticity and accuracy in generation.
@@ 3 Conclusion
Our work paves the way for several promising directions for future research. Firstly, expanding the diversity and quantity of molecular spectra data could further boost the performance of spectral-conditioned generation. Secondly, extending the framework to handle even more complex spectra modalities, such as NMR or mass spectra, could generalize the method toward broader applications in analytical chemistry. Lastly, adapting DiffSpectra to larger biomolecular or materials systems may open opportunities for molecular identification in drug discovery and materials science, especially when combined with high-throughput experimental pipelines. We believe DiffSpectra represents a significant step toward trustworthy and accurate molecular structure elucidation conditioned on spectral data.
@@ 4.1.1 Forward Diffusion Process
The SNR is strictly decreasing from $t=0$ to $t=1$, ensuring that the data is gradually corrupted into pure noise as $t\to 1$. Given two times $0\leq s<t\leq 1$, the conditional distribution of a noised graph at time $t$ given a less-noised graph at time $s$, denoted $q(\mathbf{G}_{t}\mid\mathbf{G}_{s})$, is also Gaussian:
@@ 4.1.3 Training Objective
To recover the original sample, we train a data prediction model $d_{\theta}$ that takes as input the noised graph $\mathbf{G}_{t}$, a self-conditioning estimate $\tilde{\mathbf{G}}_{0}$, and the log signal-to-noise ratio $\log\mathrm{SNR}(t)$. The model predicts the clean sample $\hat{\mathbf{G}}_{0}$, where $\hat{\mathbf{G}}_{0}$ refers to any component of $\hat{\mathcal{G}}_{0}=(\hat{\mathbf{H}},\hat{\mathbf{A}},\hat{\mathbf{X}})$. Therefore, the model is trained to minimize a weighted mean squared error:
@@ 4.2 Diffusion Molecule Transformer (DMT)
To parameterize the data prediction module $d_{\theta}$, we employ a specialized architecture named Diffusion Molecule Transformer (DMT) [37, 49] as its backbone, to jointly model the noisy molecular graph’s node features, edge features, and 3D coordinates in our diffusion framework. The architecture is motivated by the need to recover the correlations among these three components, which are independently corrupted by noise during the forward diffusion process. DMT is specifically tailored for molecular generation tasks and respects both permutation equivariance and SE(3)-equivariance.
At each time $t\in[0,1]$ during the reverse process, DMT takes the noisy molecular graph $\mathcal{G}_{t}=(\mathbf{H}_{t},\mathbf{A}_{t},\mathbf{X}_{t})$ as input, along with a self-conditioned estimate $\tilde{\mathcal{G}}_{0}$, and predicts a clean graph $\hat{\mathcal{G}}_{0}=(\hat{\mathbf{H}},\hat{\mathbf{A}},\hat{\mathbf{X}})$. The model performs denoising inference conditioned on both the current time step and the molecular spectra. The overall model architecture is illustrated in Figure 1(B). For the sake of clarity, the self-conditioning mechanism is omitted in the figure.
A time embedding, obtained via sinusoidal encoding of $\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, is passed through learnable projections. Simultaneously, a spectra embedding is generated by the SpecFormer, which will be detailed in Section 4.3. These two embeddings are then concatenated to form the conditioning vector $\mathbf{C}$, which is used throughout the network.
Initial embeddings. The DMT module is composed of stacked Transformer-style equivariant blocks. In the first layer, the node features $\mathbf{H}_{t}$ and the edge features $\mathbf{A}_{t}$ are first concatenated with their corresponding self-conditioning estimates $\tilde{\mathbf{H}}_{0}$ and $\tilde{\mathbf{A}}_{0}$, respectively. For edge features, distance features $\mathbf{D}_{0}$ computed from the self-conditioned coordinates $\tilde{\mathbf{X}}_{0}$ are also concatenated. These concatenated features are then passed through linear projections to obtain the initial embeddings:
@@
[…]
Three-stream update. Each DMT block consists of three interacting streams that are responsible for updating node features, edge features, and coordinates, respectively. These streams exchange information throughout the block. In the node stream, information is propagated across the molecular graph via a relational multi-head attention (MHA) mechanism [50, 51], enabling flexible message passing over fully connected molecular graphs. In particular, pairwise geometric distances computed from coordinates are transformed into radial basis encodings $\rho_{ij}^{(l)}$, and these are combined with edge features and geometric distances such as:
@@
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$, and $\mathbf{W}^{V}$ are learned projection matrices. The attention weights are then defined as:
@@
$\mathrm{Attn}(\mathbf{H}^{(l)},\bar{\mathbf{A}}^{(l)})_{i}=\sum_{j=1}^{N}a_{ij}\,\tanh\bigl{(}\phi_{1}(\bar{\mathbf{A}}_{ij}^{(l)})\bigr{)}\,\mathbf{V}_{j},$ (13)
with $\phi_{0}$ and $\phi_{1}$ as learnable projections. We extend the above attention mechanism to a multi-head framework, yielding the aggregated representation $\mathbf{M}^{(l)}\in\mathbb{R}^{N\times d_{m}}$, which serves as the intermediate node embeddings after relational multi-head attention.
To incorporate time-dependent and spectra-dependent conditioning, we apply adaptive layer normalization (AdaLN) to each stream, conditioned on a learned conditioning embedding $\mathbf{C}$: $\mathrm{AdaLN}({\bm{h}},\mathbf{C})=\left(1+\mathrm{FFN}_{\mathrm{scale}}(\mathbf{C})\right)\cdot\mathrm{LN}({\bm{h}})+\mathrm{FFN}_{\mathrm{bias}}(\mathbf{C})$, and use an adaptive scaling function: $\mathrm{Scale}({\bm{h}},\mathbf{C})=\mathrm{FFN}_{\mathrm{scale}}^{\prime}(\mathbf{C})\cdot{\bm{h}}$ to modulate feature amplitudes. The node update is performed by:
@@ 4.3.1 Architecture of SpecFormer
For each type of spectrum, we first segment the data independently into patches and encode these patches. The patch embeddings from all spectra are then concatenated and collectively processed by a Transformer-based encoder.
Patching. Rather than encoding each frequency point individually, we divide each spectrum into several patches. This design choice offers two key benefits: (i) It allows the model to capture local semantic features—such as absorption peaks—more effectively by grouping adjacent frequency points; and (ii) it lowers the computational cost for the following Transformer layers. Formally, a spectrum ${\bm{s}}_{i}\in\mathbb{R}^{L_{i}}$ (where $i=1,\dots,|\mathcal{S}|$) is split into patches of length $P_{i}$ with stride $D_{i}$. If $0<D_{i}<P_{i}$, patches overlap by $P_{i}-D_{i}$ points; if $D_{i}=P_{i}$, patches are non-overlapping. The patching process yields a sequence ${\bm{p}}_{i}\in\mathbb{R}^{N_{i}\times P_{i}}$, where $N_{i}=\left\lfloor\frac{L_{i}-P_{i}}{D_{i}}\right\rfloor+1$ is the number of patches produced from ${\bm{s}}_{i}$.
Patch and position encoding. Each patch sequence from the $i$-th spectrum is projected into a latent space of dimension $d$ via a learnable linear transformation $\mathbf{W}_{i}\in\mathbb{R}^{P_{i}\times d}$. To preserve the sequential order, a learnable position encoding $\mathbf{W}_{i}^{\text{pos}}\in\mathbb{R}^{N_{i}\times d}$ is added: ${\bm{p}}_{i}^{\prime}={\bm{p}}_{i}\mathbf{W}_{i}+\mathbf{W}_{i}^{\text{pos}}$, yielding the encoded representation ${\bm{p}}_{i}^{\prime}\in\mathbb{R}^{N_{i}\times d}$ for each spectrum. These representations serve as input to the SpecFormer encoder.
SpecFormer: multi-spectrum Transformer encoder. Existing encoders such as CNN-AM [52] utilize one-dimensional convolutions and are typically designed to process a single type of spectrum. In contrast, our model simultaneously considers multiple molecular spectra (e.g., UV-Vis, IR, Raman) as input. This multi-spectrum strategy is motivated by the presence of both intra-spectrum dependencies—relationships among peaks within the same spectrum—and inter-spectrum dependencies—correlations between peaks across different spectral modalities. These dependencies have been well documented, for example, in studies of vibronic coupling [53].

### v2

@@ 2.10 Elucidation trajectory analysis based on gradient
Temperature Acc@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{CosSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{TaniSim}_{\mathrm{MA}}$ $\uparrow$ FraggleSim $\uparrow$ FGSim $\uparrow$ $\tau$=0.0 14.65% 4.6948 0.4408 0.5595 0.6754 0.8136 0.7819 $\tau$=0.2 33.23% 1.9900 0.7005 0.7774 0.8763 0.9297 0.9354 $\tau$=0.4 38.31% 1.4846 0.7659 0.8287 0.9151 0.9449 0.9595 $\tau$=0.6 40.52% 1.3694 0.7826 0.8414 0.9232 0.9500 0.9637 $\tau$=0.8 40.40% 1.3190 0.7859 0.8438 0.9250 0.9512 0.9651 $\tau$=1.0 40.76% 1.3337 0.7875 0.8450 0.9262 0.9504 0.9652 $\tau$=1.2 39.59% 1.3850 0.7811 0.8402 0.9224 0.9485 0.9628
@@ 2.11 Sampling temperature balances diversity and accuracy
Since each molecular spectrum uniquely determines a molecular structure, accurate molecular structure elucidation is paramount. However, diffusion models inherently introduce stochasticity during sampling through the injection of random noise, which is essential for promoting diversity in the generated results. To better understand the trade-off between stochasticity and accuracy, we examined how modulating this randomness affects performance. Specifically, we introduce a sampling temperature parameter $\tau$ in Eq. 6, which scales the injected noise during sampling. Lower values of $\tau$ diminish stochasticity, resulting in more deterministic outputs, while higher values of $\tau$ allow for broader exploration of the molecular space.
@@
As shown in Table 4, we evaluated DiffSpectra under varying $\tau$ values ranging from 0.0 to 1.2. Moderate temperature values ($\tau=0.8$ and $\tau=1.0$) achieve the best balance between diversity and accuracy. In particular, $\tau=1.0$ obtains the highest top-1 accuracy of 40.76%, while maintaining competitive similarity metrics, including a TanimotoMG of 0.7875, CosineMG of 0.8450, and TanimotoMA of 0.9262. Extremely low temperatures (e.g., $\tau=0.0$) result in overly deterministic results with reduced accuracy and elevated MCES, while extremely high temperatures (e.g., $\tau=1.2$) also degrade performance due to excessive randomness.
@@
Taken together, these results suggest that a moderate choice of $\tau$ helps DiffSpectra strike a balance between structural diversity and high fidelity to the ground-truth molecules.
3 Discussion
In this work, we introduced DiffSpectra, the first spectra-conditioned diffusion framework for molecular structure elucidation that directly integrates multi-modal spectral information into both 2D topological and 3D geometric structure generation. Our results demonstrate that the framework not only ensures chemical validity and stability that are comparable to the training distribution but also achieves a top-1 exact accuracy of 40.8%, with functional group recovery exceeding 96%. Furthermore, by sampling multiple candidates, DiffSpectra consistently ranks the correct structure among the top 10 predictions, yielding a top-10 accuracy of 99.5%. This highlights the practical utility of diffusion-based sampling, whereby a shortlist of candidates can be reliably provided for downstream validation.
Several core findings underpin the effectiveness of DiffSpectra. First, the incorporation of 3D geometries alongside topological graphs markedly improves structure elucidation performance, underscoring the fact that spectral signatures are inherently tied to molecular geometry. Second, our comparison of SE(3) equivariance strategies shows that explicitly embedding geometric inductive biases into the model architecture provides more robust performance than relying solely on data augmentation, emphasizing the importance of principled symmetry constraints. Third, conditioning on multiple spectral modalities leads to complementary structural information: while Raman spectra provide richer constraints than UV-Vis, the integration of Raman, IR, and UV-Vis yields the most accurate elucidations. Fourth, pre-training the spectral encoder SpecFormer with masked reconstruction and contrastive alignment proves highly beneficial, providing inductive biases that significantly improve spectral–structural alignment. Finally, gradient-based trajectory analysis reveals a mechanistic view of the generative process: DiffSpectra first establishes global semantic information across atoms, then elucidates simple substructures, subsequently resolves more complex functional groups, and ultimately converges to a chemically consistent molecular structure. This staged generative trajectory mirrors the reasoning process of human chemists and provides interpretability for the model’s predictions.
The demonstrated performance and interpretability of DiffSpectra have direct implications for practical applications in natural product discovery, drug development, and functional materials design. By providing high-quality candidate structures directly from spectra, DiffSpectra offers a scalable alternative to traditional manual elucidation pipelines and opens opportunities for integrating automated generative models into real-world experimental workflows.
Despite its encouraging performance, DiffSpectra currently focuses on three spectral types—UV-Vis, IR, and Raman—so its applicability to other modalities, such as nuclear magnetic resonance (NMR) or mass spectrometry (MS), remains to be further explored. In addition, our validation to date has been primarily computational. Therefore, extensive experimental studies will be an important next step toward further strengthening the robustness of the framework in practical laboratory settings.
@@ 3 Discussion
Our work paves the way for several promising research directions. Expanding the diversity and scale of available spectral datasets will likely further improve the generalizability and performance of spectra-conditioned generation. Extending the framework to handle more complex modalities such as NMR, MS, or X-ray spectra would enhance its applicability across a wider range of analytical chemistry tasks. Moreover, adapting DiffSpectra to larger biomolecular and materials systems could unlock new opportunities for molecular identification in drug discovery and materials science, particularly when coupled with high-throughput experimental pipelines.
In summary, DiffSpectra represents a significant advance in molecular structure elucidation by unifying topological, geometric, and spectral reasoning within a diffusion-based generative framework. By demonstrating high accuracy, interpretability, and generalizability, DiffSpectra establishes a new paradigm for automated and scalable structure elucidation. We believe that this work makes an important step toward trustworthy and accurate spectra-conditioned molecular structure elucidation, with broad potential to impact computational chemistry, drug discovery, and materials science.
@@ 4.1.1 Forward Diffusion Process
The SNR is strictly decreasing from $t=0$ to $t=1$, ensuring that the data is gradually corrupted into pure noise as $t\to 1$. Given two time steps, $s$ and $t$, $0\leq s<t\leq 1$, the conditional distribution of a noised graph at time $t$ given a less-noised graph at time $s$, denoted $q(\mathbf{G}_{t}\mid\mathbf{G}_{s})$, is also Gaussian:
@@ 4.1.3 Training Objective
[…]
@@ 4.2 Diffusion Molecule Transformer (DMT)
To parameterize the data prediction module $d_{\theta}$, we employ a specialized architecture, the Diffusion Molecule Transformer (DMT) [45, 51], as its backbone, to jointly model the node features, edge features, and 3D coordinates of the noisy molecular graph within our diffusion framework. The architecture is motivated by the need to recover the correlations among these three components, each of which is independently corrupted by noise during the forward diffusion process. DMT is specifically tailored for molecular generation tasks and preserves both permutation equivariance and SE(3)-equivariance.
[…]
A time embedding, obtained via sinusoidal encoding of $\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, is passed through learnable projections. Simultaneously, embeddings of spectra are generated by the SpecFormer, which will be detailed in Section 4.3. These two embeddings are then concatenated to form the conditioning vector $\mathbf{C}$, which is used throughout the network.
Initial embeddings. The DMT module consists of stacked Transformer-style equivariant blocks. In the first layer, the node features $\mathbf{H}_{t}$ and the edge features $\mathbf{A}_{t}$ are concatenated with their corresponding self-conditioning estimates $\tilde{\mathbf{H}}_{0}$ and $\tilde{\mathbf{A}}_{0}$, respectively. For edge features, distance features $\mathbf{D}_{0}$ computed from the self-conditioned coordinates $\tilde{\mathbf{X}}_{0}$ are also concatenated. These concatenated features are then passed through linear projections to obtain the initial embeddings:
@@
where $\left[\cdot,\cdot\right]$ denotes the concatenation operation, and $d_{h}$ and $d_{a}$ denote the hidden embedding dimensions for node and edge embeddings, respectively. The coordinate stream initializes with the noisy positions $\mathbf{X}^{(1)}=\mathbf{X}_{t}$. In addition, adjacency matrices derived from chemical bonding patterns and distance cutoffs computed from $\tilde{\mathbf{X}}_{0}$ are extracted and concatenated to serve as auxiliary attention masks in subsequent layers.
Three-stream update. Each DMT block consists of three interacting streams responsible for updating node features, edge features, and coordinates, respectively. These streams exchange information throughout the block. In the node stream, information is propagated across the molecular graph via a relational multi-head attention (MHA) mechanism [52, 53], enabling flexible message passing over fully connected molecular graphs. In particular, pairwise geometric distances computed from coordinates are transformed into radial basis encodings $\rho_{ij}^{(l)}$, and these are combined with edge features and geometric distances such as:
@@
where $\mathbf{W}^{Q},\mathbf{W}^{K},\mathbf{W}^{V}$ are learned projection matrices. The attention weights are then defined as:
@@
$\mathrm{Attention}(\mathbf{H}^{(l)},\bar{\mathbf{A}}^{(l)})_{i}=\sum_{j=1}^{N}a_{ij}\,\tanh\bigl(\phi_{1}(\bar{\mathbf{A}}_{ij}^{(l)})\bigr)\,\mathbf{V}_{j},$ (13)
with $\phi_{0}(\cdot)$ and $\phi_{1}(\cdot)$ as learnable projections. We extend the above attention mechanism to a multi-head relational attention framework. For each head $h=1,\dots,H$, we compute head-specific queries, keys, and values as:
$\mathbf{Q}_{h}=\mathbf{H}^{(l)}\mathbf{W}_{h}^{Q},\quad\mathbf{K}_{h}=\mathbf{H}^{(l)}\mathbf{W}_{h}^{K},\quad\mathbf{V}_{h}=\mathbf{H}^{(l)}\mathbf{W}_{h}^{V},$ (14)
where $\mathbf{W}_{h}^{Q},\mathbf{W}_{h}^{K},\mathbf{W}_{h}^{V}$ are learnable projection matrices for head $h$. Each head independently performs geometry-aware attention as:
$\mathbf{M}_{h}^{(l)}=\mathrm{Attention}_{h}\bigl(\mathbf{H}^{(l)},\bar{\mathbf{A}}^{(l)}\bigr),$ (15)
where $\mathrm{Attention}_{h}(\cdot)$ denotes the relational attention operator defined above. The outputs from all heads are then concatenated and projected back to the model dimension:
$\mathbf{M}^{(l)}=\mathrm{Concat}\bigl[\mathbf{M}_{1}^{(l)},\dots,\mathbf{M}_{H}^{(l)}\bigr].$ (16)
The resulting $\mathbf{M}^{(l)}\in\mathbb{R}^{N\times d_{m}}$ serves as the aggregated node embedding enriched by geometry-aware multi-head interactions.
To incorporate time-dependent and spectra-dependent conditioning, we apply adaptive layer normalization (AdaLN) [54, 55] to each stream, conditioned on a learned conditioning embedding $\mathbf{C}$: $\mathrm{AdaLN}({\bm{h}},\mathbf{C})=\left(1+\mathrm{FFN}_{\mathrm{scale}}(\mathbf{C})\right)\cdot\mathrm{LN}({\bm{h}})+\mathrm{FFN}_{\mathrm{bias}}(\mathbf{C})$, where $\mathrm{FFN}_{\mathrm{scale}}(\cdot)$ and $\mathrm{FFN}_{\mathrm{bias}}(\cdot)$ are lightweight feed-forward networks that project the conditioning signal into scale and shift parameters. In addition, we introduce an adaptive scaling function to further modulate feature magnitude: $\mathrm{Scale}({\bm{h}},\mathbf{C})=\mathrm{FFN}_{\mathrm{scale}}^{\prime}(\mathbf{C})\cdot{\bm{h}}$, where $\mathrm{FFN}_{\mathrm{scale}}^{\prime}(\cdot)$ shares the same architecture as the above FFNs but produces a multiplicative gating factor. The node update is performed by:
@@ 4.3.1 Architecture of SpecFormer
For each type of spectrum, we first segment the data into patches and independently encode these patches. The patch embeddings from all spectra are then concatenated and collectively processed by a Transformer-based encoder.
Patching. Rather than encoding each frequency point individually, we divide each spectrum into multiple patches. This design choice offers two key benefits: (i) it enables the model to capture local semantic features—such as absorption peaks—more effectively by grouping adjacent frequency points; and (ii) it reduces the computational cost for the following Transformer layers. Formally, a spectrum ${\bm{s}}_{i}\in\mathbb{R}^{L_{i}}$ (where $i=1,\dots,|\mathcal{S}|$) is split into patches of length $P_{i}$ with stride $D_{i}$. If $0<D_{i}<P_{i}$, patches overlap by $P_{i}-D_{i}$ points; if $D_{i}=P_{i}$, patches are non-overlapping. The patching process yields a sequence ${\bm{p}}_{i}\in\mathbb{R}^{N_{i}\times P_{i}}$, where $N_{i}=\left\lfloor\frac{L_{i}-P_{i}}{D_{i}}\right\rfloor+1$ is the number of patches produced from ${\bm{s}}_{i}$.
Patch and position encoding. Each patch sequence from the $i$-th spectrum is projected into a latent space of dimension $d$ via a learnable linear transformation $\mathbf{W}_{i}\in\mathbb{R}^{P_{i}\times d}$. To preserve the sequential order, a learnable positional encoding $\mathbf{W}_{i}^{\text{pos}}\in\mathbb{R}^{N_{i}\times d}$ is added: ${\bm{p}}_{i}^{\prime}={\bm{p}}_{i}\mathbf{W}_{i}+\mathbf{W}_{i}^{\text{pos}}$, yielding the encoded representation ${\bm{p}}_{i}^{\prime}\in\mathbb{R}^{N_{i}\times d}$ for each spectrum. These representations serve as inputs to the SpecFormer encoder.
SpecFormer: multi-spectrum Transformer encoder. Existing encoders, such as CNN-AM [56], utilize one-dimensional convolutions and are typically designed to process a single type of spectrum. In contrast, our model simultaneously considers multiple molecular spectra (e.g., UV-Vis, IR, Raman) as input. This multi-spectrum strategy is motivated by the presence of both intra-spectrum dependencies—relationships among peaks within the same spectrum—and inter-spectrum dependencies—correlations between peaks across different spectral modalities. These dependencies have been well documented, for example, in studies on vibronic coupling [57].